# RefVTON: person-to-person Try on with Additional Unpaired Visual Reference

### v3

@@ 3.4 Virtual-Tryon Generation with Extra Visual Reference
For humans, particularly in e-commerce shopping, we believe that relying solely on garment images and conditions extracted by external models, as done in previous methods, is insufficient to imagine the realistic visual effects of fitting a target garment onto one’s body. A garment’s style, texture, and fine design elements are more faithfully represented when worn by another person, whereas viewing the garment in isolation fails to convey these nuances. The same applies to generative virtual try-on models, as illustrated in Fig. 2. Therefore, we argue that relying on cloth images and conditions extracted by external models, as in previous methods, is insufficient to capture the realistic visual effect of fitting a target garment onto a person. By incorporating reference person images, that the target clothes worn by a different person $\bm{r}_{i}$, into the virtual try-on generation process, the model can similarly get more direct and intuitive visual guidance.
To provide the model with reference person images $\bm{r}_{i}$, we construct pairs in the form of “different persons wearing the target garment,” denoted as $\bm{c}_{i},\bm{r}_{i}]$. Such reference images $\bm{r}_{i}$ are not available in existing open-source virtual try-on datasets, similar to the person images $\bm{p}_{i}$. In this subsection, we introduce our reference data generation pipeline, which synthesizes the required reference images, enriching current virtual try-on datasets with supplementary references and enabling our model to be trained with additional visual guidance.
@@
Existing open-source datasets, including VITON-HD [choi2021viton], DressCode [morelli2022dress], and IGPairs [shen2025imagdressing], lack unpaired reference images $\bm{r}_{i}$ showing the different persons wearing the target garment, thereby constraining the model’s training. We employ generative models to synthesize such reference images to overcome this limitation. For the generated reference data $\bm{r}_{i}$ to provide accurate and intuitive visual guidance, we argue that they should meet the following requirements:
@@
Preserve the target garment faithfully. The target clothing’s color, texture, and design must remain unchanged to ensure accurate reference.
@@
Introduce diversity in person’s appearance. The person wearing the target cloth in the reference image should unlike the target person wearing the same garment. Otherwise, the model may learn shortcut and overfit to the target image. This diversity can be achieved by altering hairstyle, hair/skin color, body pose, or facial expressions.
@@
Vary the non-target garments to provide outfit diversity. While the target garment remains unchanged, other garments should be modified. For example, if the target garment is an upper-body item, the reference image should retain the same upper-body garment but alter the lower-body clothing, shoes, or accessories.
The overall generation pipeline for reference data is illustrated in Figure 5.
As shown in the pipeline, we employ Flux-Kontext [labs2025flux1kontextflowmatching] to generate reference images from the target descriptions, leveraging its strong capability to maintain consistency with the input image (1). To address requirement (2), we describe the target image and then intentionally provide an alternative description that ensures the reference image differs from the target input. For this purpose, we use Qwen2.5-VL [bai2025qwen2] to generate detailed descriptions of the model’s appearance, and instruct it to produce variants that do not resemble the original model. Finally, to fulfill requirement (3), we curate a list of garment descriptions representing diverse non-target clothing items. The outlook, action, and outfit descriptions are concatenated as the positive prompt. In contrast, the original image description is used as the negative prompt, and both are fed into the generative model to synthesize reference images.
Figure 6: Qualitative comparison on the VITON dataset. and the model is trained following the pipeline in Fig. 3 (b). “Ref” denotes the additional reference $\bm{r}_{i}$ image is used during the inference, while “mask-free” indicates that the image is generated using an unmasked person $\bm{p}_{i}$ image instead of a masked agnostic image $\bm{a}_{i}$.
@@
We supplement existing virtual try-on benchmarks including VITON-HD [han2018viton], DressCode [morelli2022dress], ViViD [fang2024vivid], FashionTryOn [zheng2019virtually] and IGPairs [shen2025imagdressing] by generating corresponding reference pairs for each target garment image $\bm{c}_{i}$, forming data pairs $\bm{c}_{i},\bm{r}_{i}]$ that are used for both model training and evaluation.Some open-source datasets, such as FashionTryOn [zheng2019virtually] and IGPairs [shen2025imagdressing], contain numerous duplicated or low-quality samples. We compare the CLIP features of images to filter out redundant samples and employ Qwen2.5-VL to identify distorted or unclear images, as well as images where the person faces away from the camera, ensuring overall data quality before generating the final reference set via our data generation pipeline. Finally, we enrich open-source virtual try-on datasets with additional visual references $\bm{r}_{i}$ and person images $\bm{p}_{i}$, and combine them to form our own dataset, named Virtual Fitting with Reference (VFR), for training our RefTON model. The detailed data collection, filtering procedures, and sample’s visualization are provided in the Appendix.
@@ 4 Experiments
We evaluate the effectiveness of our proposed method on two public benchmarks, DressCode [morelli2022dress] and VITON-HD [choi2021viton], both containing images with a resolution of $1024\times 768$. The VITON-HD dataset comprises 13,670 upper-body image pairs of women, split into 11,647 pairs for training and 2,032 pairs for testing. The DressCode dataset includes three subsets—upper body, lower body, and dresses—with 48,392 training and 5,400 testing pairs. Since DressCode does not provide wrapped cloth masks or agnostic images $\bm{a}_{i}$, we generate them using the mask generation tool from CatVTON [chong2024catvton]. During training, we fine-tune our model on the flux-kontext backbone using the Low-Rank Adaptation (LoRA) technique, with a rank of $64$ and $\alpha=128$, optimized by AdamW. For quantitative evaluation, all generated images are resized to $512\times 384$ to ensure a fair comparison with previous methods. In a single-dataset experiment, the model is trained independently for 20,000 steps on VITON-HD and 48,000 on DressCode with a batch size of 128, using 8 NVIDIA H100 GPUs. To further enhance generalization and robustness to in-the-wild inputs, we train an additional model on the mixed VFR dataset introduced in Sec. 3.4. We report cross-dataset evaluation results of our RefVTON model at a resolution of $1024\times 768$ on both VITON-HD and DressCode, and demonstrate its in-the-wild performance using images captured by ourselves and collected from the internet.
@@ 4.1 Quantative result
We evaluate the numerical results of our virtual try-on model on the VITON and DressCode datasets, distinguishing between paired and unpaired try-on settings. For paired try-on settings with ground truth in test datasets, we utilize four widely used metrics to assess the similarity between the synthesized images and their corresponding authentic images: the Structural Similarity Index (SSIM) [nilsson2020understanding], Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable], Fréchet Inception Distance (FID)[Seitzer2020FID], and Kernel Inception Distance (KID)[binkowski2018demystifying]. For unpaired settings, where we measure the distributional similarity between the synthesized and real samples, we specifically rely on the Fréchet Inception Distance (FID) and Kernel Inception Distance (KID). As shown in Table4.
Our method (RefVTON) consistently performs better than prior baselines, demonstrating higher try-on fidelity and strong alignment with the target person’s pose. With the addition of reference images (“+R”), the quality and detail consistency of the try-on results is further improved compared with the results without reference images, establishing new state-of-the-art results across multiple metrics. Notably, even in the mask-free setting—without agnostic masks or auxiliary inputs—our method maintains garment style correctness and pose consistency, while reaching accuracy on par with or superior to baseline methods, highlighting its robustness and practicality.
Table 1: Quantitative comparison across VITON-HD [choi2021viton] and DressCode [morelli2022dress]. The best and second best results are shown in bold and underline. “+R” denotes the use of reference images, and “MF” indicates mask-free inputs. Missing values are shown as “–”. Subscripts $p$ and $u$ represent the paired and unpaired test settings, respectively. Unless otherwise specified, the same notations carry the same meanings throughout the figures and tables in this paper.
@@
CatVTON(MF) [chong2024catvton] 0.061 0.870 5.89 0.51 9.29 1.17 0.045 0.902 4.78 1.30 7.40 2.62

### v5

@@ 3.4 Virtual-Tryon Generation with Extra Visual Reference
For both humans and generative virtual try-on models, garment images and conditions extracted by external models alone are insufficient to capture the realistic visual effect of wearing a target garment, as illustrated in Fig. 2. A garment’s style, texture, and fine design details are more faithfully presented when worn by another person than when shown in isolation. Therefore, we introduce reference person images $\bm{r}_{i}$, in which another person wears the target garment, to provide more intuitive visual guidance during virtual try-on generation. To obtain such references, we construct pairs of the form $\bm{c}_{i},\bm{r}_{i}]$, representing ”different persons wearing the target garment.” Since these reference images $\bm{r}_{i}$ are unavailable in existing open-source virtual try-on datasets, we develop a reference data generation pipeline to synthesize them. This pipeline augments current datasets with supplementary references and enables model training with additional visual guidance.
@@
Existing open-source datasets like VITON-HD [choi2021viton], DressCode [morelli2022dress], and IGPairs [shen2025imagdressing], lack unpaired reference images $\bm{r}_{i}$ showing different persons wearing the target garment. To overcome this limitation, we employ editing models to synthesize such reference images. To serve as accurate and informative visual guidance, the generated reference data $\bm{r}_{i}$ should satisfy the following requirements:
@@
  • Preserve the target garment faithfully. The target garment’s color, texture, and design must remain unchanged to ensure an accurate reference.
@@
  • Introduce diversity in the person’s appearance. The person wearing the target garment in the reference image should be different from the target person wearing the same garment. Otherwise, the model may learn shortcuts and overfit to the target image. This diversity can be achieved by altering hairstyle, hair color, skin tone, body pose, or facial expression.
@@
  • Vary the non-target garments to provide outfit diversity. While the target garment remains unchanged, other garments should be modified. For example, if the target garment is an upper-body item, the reference image should retain the same upper-body garment while altering the lower-body clothing, shoes, or accessories.
Since the reference image is edited from the target image $\bm{p}_{i,\bm{c}_{i}}$ in existing datasets, the first requirement ensures faithful garment preservation, while the second and third promote diversity in non-target regions to prevent overfitting and prevent the model from taking the shortcut of directly copying the target image in the generated try-on results.
As illustrated in the pipeline, we use Flux-kontext [labs2025flux1kontextflowmatching] to synthesize reference-person images. Given a target image and carefully designed text prompts, Flux-kontext generates images that faithfully preserve the target garment’s color, texture, and design while varying the wearer’s appearance and non-target clothing. Specifically, we first feed the target image and a textual instruction into Qwen2.5-VL [bai2025qwen2] to obtain a detailed description of the person’s appearance, together with an opposite description. We then collect a set of non-target garments and action descriptions to introduce diversity in clothing and pose; examples of these garments and action items are provided in Appendix A. Finally, we concatenate the opposite appearance description with the action and non-target garment descriptions to form the positive prompt, while the original appearance description is used as the negative prompt. These prompts are then fed into Flux-kontext to synthesize the reference images.
@@
We supplement existing virtual try-on benchmarks including VITON-HD [han2018viton], DressCode [morelli2022dress], ViViD [fang2024vivid], FashionTryOn [zheng2019virtually] and IGPairs [shen2025imagdressing] by generating corresponding reference pairs for each target garment image $\bm{c}_{i}$, forming data pairs $\bm{c}_{i},\bm{r}_{i}]$ that are used for both model training and evaluation.Some open-source datasets, such as FashionTryOn [zheng2019virtually] and IGPairs [shen2025imagdressing], contain numerous duplicated or low-quality samples. We compare the CLIP features of images to filter out redundant samples and employ Qwen2.5-VL to identify distorted or unclear images, as well as images where the person faces away from the camera, ensuring overall data quality before generating the final reference set via our data generation pipeline. Finally, we enrich open-source virtual try-on datasets with additional visual references $\bm{r}_{i}$ and person images $\bm{\bar{p}}_{i,c_{j}}$, and combine them to form our own dataset, named Virtual Fitting with Reference (VFR), for training our RefTon model. The detailed data collection, filtering procedures, and visualization of the samples are provided in Appendix A.
Figure 6: Qualitative comparison on the VITON-HD dataset. “reference” denotes using additional reference $\bm{r}_{i}$ for the inference.
Figure 7: Qualitative comparison on the DressCode dataset., and the model is trained following the pipeline in Fig. 3 (b). “reference” denotes the reference $\bm{r}_{i}$ image is used during the inference.
@@ 4 Experiments
We mainly evaluate our method on two public benchmarks, DressCode [morelli2022dress] and VITON-HD [choi2021viton], both containing images with a resolution of $1024\times 768$. The VITON-HD comprises 13,670 upper-body image pairs of women, split into 11,647 pairs for training and 2,032 pairs for testing. The DressCode includes three subsets—upper body, lower body, and dresses—with 48,392 training and 5,400 testing pairs. Since DressCode does not provide wrapped cloth masks or agnostic images $\bm{a}_{i}$, we generate them using the mask generation tool from CatVTON [chong2024catvton]. We further conducted experiments on the StreetTryOn [cui2023street-tryon] and included comparisons with the methods mentioned to demonstrate the model’s P2P and in-the-wild capabilities. In the P2P evaluation, we did not use the cloth-only image as input and only used the person image as a reference.
During training, we fine-tune our model on the flux-kontext backbone using the Low-Rank Adaptation (LoRA) technique, with a rank of $64$ and $\alpha=128$, optimized by AdamW. For quantitative evaluation, all generated images are resized to $512\times 384$ to ensure a fair comparison with previous methods. In a single-dataset experiment, the model is trained independently for 20,000 steps on VITON-HD and 48,000 on DressCode with a batch size of 128, using 8 NVIDIA H100 GPUs. To further enhance generalization and robustness to in-the-wild inputs, we train an additional model on the mixed VFR dataset introduced in Sec. 3.4. We report cross-dataset evaluation results of our RefTon model at a resolution of $1024\times 768$ on both VITON-HD and DressCode, and demonstrate its in-the-wild performance using images collected by ourselves.
@@ 4.1 Quantative result
We provide the numerical results of our model on the VITON-HD and DressCode datasets, distinguishing between paired and unpaired try-on settings. For paired try-on settings with ground truth in test datasets, we utilize four metrics to assess the similarity between the synthesized images and their corresponding authentic images: the Structural Similarity Index (SSIM) [nilsson2020understanding], Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable], Fréchet Inception Distance (FID)[Seitzer2020FID], and Kernel Inception Distance (KID)[binkowski2018demystifying]. For unpaired settings, where we measure the distributional similarity between the synthesized and real samples, we specifically rely on the Fréchet Inception Distance (FID) and Kernel Inception Distance (KID). As shown in Table 1.
Table 1: Quantitative comparison on VITON-HD [choi2021viton] and DressCode [morelli2022dress]. Best and second-best results are shown in bold and underline, respectively. “+R” denotes the use of reference images, “MF” indicates mask-free inputs, and “–” denotes missing values. Subscripts $p$ and $u$ denote the paired and unpaired settings, respectively.
@@
CatVTON(Mask-Free) [chong2024catvton] 0.061 0.870 5.89 0.51 9.29 1.17 0.045 0.902 4.78 1.30 7.40 2.62
@@
Our method (RefTon) consistently performs better than prior baselines, demonstrating higher try-on fidelity and strong alignment with the target person’s pose. With the addition of reference images (“+R”), the quality and detail consistency of the try-on results are further improved compared with the results without reference images, establishing new state-of-the-art results across multiple metrics. Notably, even in the mask-free setting—without agnostic masks or auxiliary inputs—our method maintains garment style correctness and pose consistency, while reaching accuracy on par with or superior to baseline methods, highlighting its robustness and practicality.
Table 1 summarizes the quantitative evaluation on the DressCode dataset. Our method (RefTon) outperforms all baselines, delivering higher try-on quality and achieving strong consistency with the target person’s pose and body structure. Integrating reference images (“+R”) further enhances the results, establishing a new state-of-the-art. Importantly, even in the mask-free setting—without agnostic masks or additional inputs—our method correctly preserves garment styles (e.g., clothing length and design) and maintains high pose alignment, while achieving accuracy comparable to or surpassing prior baselines, demonstrating robustness and practicality. Additional quantitative results on the DressCode subset are provided in Appendix B in detail.
Results in Table 2 show that our method achieves state-of-the-art performance on the StreetTryOn benchmark. Notably, our model is neither trained on the StreetTryOn domain nor designed to handle the absence of garment images during training. These results demonstrate its strong generalization ability and flexibility, and further confirm that it can directly extract garment information from the reference image for virtual try-on.
Table 2: Quantitative comparisons on the StreetTryOn. Sh, St, P, denotes the shop, street, and person(model) image, respectively.
Method Required Input Sh-to-St P-to-P P-to-St St-to-St
Mask Pose Text $\downarrow FID\downarrow$ $\downarrow FID\downarrow$
StreetTryOn 34.054 12.185 34.191 33.039
OmniVTON 33.919 8.983 33.450 23.470
RefTON 28.991 8.870 25.429 16.452
4.2 Qualitative comparison
Previous methods can fit garments onto a person, but often fail to preserve wearing-critical attributes such as cut, style, and fine details. By contrast, our method achieves improved visual fidelity among all baselines. As shown in Fig. 6, on VITON, RefTon realistically renders challenging materials such as hollow and semi-transparent fabrics even without reference images. For example, it faithfully preserves the perforated structures and transparency of lace garments, while baselines often produce solid textures or spurious dotted artifacts. Our method also better maintains garment patterns, keeping printed letters and logos clear and consistent with the input. Adding reference images further improves generation quality, validating the effectiveness of our reference assist framework. Even without agnostic masks, directly transferring clothing $\bm{c}_{i}$ onto person images $\bm{\bar{p}}_{i,\bm{c}_{j}}$ still allows our method to outperform most baselines. More results on VITON-HD, DressCode, and in-the-wild examples are provided in Appendix C.
Table 3: Cross-dataset comparison with OOTDiffusion [xu2025ootdiffusion] on VITON-HD and DressCode under both paired and unpaired settings. For the mask-free setting, we report only the unpaired results, as the paired setting is not meaningful when no garment–person alignment is enforced.
Methods VITON-HD (Paired) DressCode (Paired)
[…]
RefTON (Ours) 0.851 6.23 0.80 0.072 0.896 3.70 1.13 0.045
RefTON+R (Ours) 0.859 5.13 0.62 0.060 0.903 3.14 0.97 0.038
Methods VITON-HD (Unpaired) DressCode (Unpaired)
$downarrow FIDu\downarrow KIDu\downarrow$
RefTON (Ours) 9.11 1.08 5.22 1.20
RefTON+R (Ours) 8.59 0.87 5.03 1.11
RefTON/MF (Ours) 8.88 0.82 5.03 1.23
RefTON+R/MF (Ours) 8.39 0.65 4.87 1.10
Our model also performs well on a benchmark with a wider variety of clothing types. As shown in Fig. 7, we evaluate our method on the DressCode dataset, where garments are categorized into upper body, lower body, and dresses. Our approach produces more faithful and natural try-on results compared to previous methods. In particular, it renders reflective materials, such as leather and metallic fabrics, with superior realism, avoiding the over-smoothing or distortion artifacts commonly observed in other methods. Moreover, even without agnostic masks, our model can still perform consistent try-on guided by garment style, accurately preserving the length, structure, and overall design without introducing mismatched or inconsistent shapes.
4.3 Training on VRF Dataset and Evaluation
To evaluate the effectiveness of our person-to-person virtual try-on framework and data construction pipeline, we build a mixed dataset, Mixed-Virtual-Ref, by combining samples from DressCode, VITON-HD, FashionTryOn [xiao2025omnigen], ViViD [fang2024vivid], and IGPairs [choi2024improving]. We use Qwen2.5-VL to filter low-quality images, generate agnostic masks via [chong2024catvton], and obtain reference and unpaired person images through our data generation process. In total, 103,936 image pairs are collected for training, using the same hyperparameters as in the DressCode and VITON-HD experiments.
We evaluate the model on the DressCode and VITON-HD test sets, with quantitative results shown in Table 3. Our method (RefTon) achieves consistently superior or comparable performance to OOTDiffusion [xu2025ootdiffusion] across both benchmarks and both paired and unpaired settings. Notably, despite not being trained on DressCode or VITON-HD individually, the mixed-dataset model surpasses dataset-specific baselines on most metrics (e.g., FID and LPIPS), demonstrating strong cross-dataset generalization and the robustness of our data construction strategy.
4.4 Ablation Study
To validate the rescaled position index, we experiment on VITON-HD with cloth image $\bm{c}_{i}$, dense pose $\bm{d}_{i}$, warp mask $\bm{m}_{i}$, and reference image $\bm{r}_{i}$ as conditions. In the masked and mask-free settings, the target cloth is transferred to the person image and the agnostic image, respectively. We evaluate two condition scales, $0.25\times$ and $0.5\times$, where the person or agnostic image is resized to the target resolution and the conditional inputs are rescaled accordingly. As shown in Table 4, the rescaled position index outperforms the original Flux-Kontext position index on FID and KID under both settings, showing its effectiveness for multi-condition inputs with varying resolutions.
Table 4: Rescaled position index (PI) ablation.. We compare the original Flux-Kontext position index with our rescaled PI at different condition scales; MFMF denotes mask-free inputs. Conditional inputs are resized to $0.5\times$ and $0.25\times$ of the target resolution.
Method 0.5 Scale 0.25 Scale
[…]
w/ o Rescaled PI 5.29 0.84 4.75 0.72 6.08 0.93 5.35 0.76
w/ Rescaled PI 5.09 0.79 4.71 0.69 6.01 0.77 5.37 0.71
5 Conclusion
This paper introduces RefTon, a virtual try-on framework that supports both mask-based and mask-free inference and leverages reference images to guide the try-on process. We extend Flux-Kontext to handle multi-condition inputs of varying resolutions via a rescaled position index. To train RefTon, we propose a reference data generation pipeline integrating Qwen2.5-VL and Flux-Kontext. This design allows RefTon to faithfully preserve translucent fabrics, intricate designs, and fine details, consistently outperforming existing methods both quantitatively and qualitatively, achieving state-of-the-art performance in all settings.
6 Acknowledgments
This research was supported by 360 AI Research. We sincerely thank 360 AI Research for its support of this work.
Supplementary Material
Appendix A Generation of Reference Image
Firstly, the reference image should faithfully preserve the details of the target garment $\bm{c}_{i}$. This requires the editing model to have strong consistency abilities. The Flux-Kontext model has a robust ability to edit the target region corresponding to the text prompt while keeping unrelated regions—such as the area not related to the garment—unchanged. Moreover, the Flux-Kontext model can perform precise local editing according to the text prompt, in this case, focusing on the person under the garment. Therefore, we choose the popular Flux-Kontext model [36] to edit the input image conditioned on the text prompt. Specifically, we add a sentence such as “keep the {target cloth} cloth unchanged” in the text prompt.
Secondly, the person in the reference image should look significantly different from the person in the target image. During training, we observed that if the reference image is too similar to the target image, the model tends to rely on a ”shortcut”—directly copying from the reference image and ignoring the agnostic/person image $\bm{a}_{i}$$\bm{p}_{i}$ and the cloth $\bm{c}_{i}$. To avoid this, we ensure that the person in the reference image differs from the target image to better showcase the visual effect of the clothing when worn, rather than focusing on the appearance of the person themselves. To achieve this, we utilize the Text-to-Image (T2I) capabilities of the Flux-Kontext model. We extract an accurate description of the person’s appearance in the target image (e.g., ”The model has an East Asian appearance, with light skin, long black hair, and a neutral expression…”) and pass it as the negative prompt to the T2I model. In contrast, we provide an opposite description (e.g., ”The model has an African appearance, with dark skin, short yellow hair, and a cheerful expression…”) as the positive prompt to guide the editing of the target image, as shown in Figure 4(b).
Fig S9 illustrates selected text prompts from the prompt description bank used for reference image generation. Furthermore, Fig S10 exhibits a selection of the resulting reference data samples.
Figure S9: Sample text prompts from the Outfit and Action Description Bank. To ensure the model edits only the person while preserving the target clothing, we assign different outfits and action description categories to different clothing inputs.
Figure S10: Sample reference images generated by our reference data generation pipeline. The editing model takes the target person’s image as input and synthesizes corresponding reference images, while preserving the garment’s appearance to match the cloth.
Appendix B Further Quantitative Evaluation on High Resolution
Here we provide more results of the generated image. Table S5 summarizes the detailed quantitative evaluation on the DressCode dataset. Our method (RefTON) outperforms all baselines, delivering higher try-on quality and achieving strong consistency with the target person’s pose and body structure. Integrating reference images (“+R”) further enhances the results, establishing a new state-of-the-art. Importantly, even in the mask-free setting—without agnostic masks or additional inputs—our method correctly preserves garment styles (e.g., clothing length and design) and maintains high pose alignment, while achieving accuracy comparable to or surpassing prior baselines, demonstrating robustness and practicality.